# On the Complexity of Determinations

### 30 Mar 2026 v1

@@ Abstract
Classical complexity theory measures the cost of computing a function, but many computational tasks require committing to one valid output among several. We introduce determination depth—the minimum number of sequential layers of irrevocable commitments needed to select a single valid output—and show it is an orthogonal axis measuring the cost of committing. We exhibit relational tasks whose commitments are constant-time table lookups yet require exponential parallel width to compensate for any reduction in depth. A conservation law couples determination and computational depth: enriching the commitment basis trades layers for step complexity, but the total sequential cost is bounded below. For circuit-encoded specifications, the resulting depth hierarchy captures the polynomial hierarchy ($\Sigma_{2k}^{P}$-complete for each fixed $k$, PSPACE-complete for unbounded $k$). Determination depth and computational depth are orthogonal: neither subsumes the other. Stable matching witnesses the independence sharply—finding a stable matching is in P, yet every finite determination depth arises as the rotation-poset height of some instance. In the distributed setting, the framework recovers the Halpern–Moses common-knowledge impossibility.
@@ 1 Introduction
Classical complexity theory measures the cost of computing a function—producing the unique correct output. But many tasks require resolving a relation: choosing one admissible outcome from a set, as when a text generator selects one coherent continuation among many, or a distributed system must agree on one valid output among several. We show that this distinction creates a new, orthogonal axis of complexity. Classical hard functions are computationally expensive but have no cost on this new axis. However, there exist relational tasks in which every step is a constant-time table lookup, yet any strategy using fewer than $k$ sequential layers requires exponential width (exponentially many parallel candidates), even with unbounded parallelism within each candidate. The bottleneck is not computation, nor communication, but irrevocable commitment. This cost is invisible to classical complexity measures—circuit depth, communication complexity, adaptive queries—because they track the cost of computing or distributing an answer, not the cost of committing to one.
@@
We establish three main results. (1) Exponential separation (Section 3): the depth–width tradeoff is exponential, generalizing pointer chasing [NW93, MYZ25] from functions to relations. (2) Conservation law (Section 4): enriching the commitment basis trades determination layers for computational depth within layers, but the total sequential cost is bounded below—choices can be bundled into richer steps, but the burden does not disappear. The bound is tight. Under a commutative basis in the online setting, a supporting oracle characterization shows that determination depth equals the minimum number of observations of the growing history, regardless of computational power between observations. (3) Polynomial-hierarchy characterization (Section 5): for circuit-encoded specifications, “is depth $\leq k$?” is $\Sigma_{2k}^{P}$-complete for each fixed $k$ and PSPACE-complete for unbounded $k$, certifying determination depth as a semantic measure well-calibrated against the classical hierarchy.
@@
The appendices exhibit both phenomena. BSP round complexity overcharges by counting communication rounds that involve no commitment, and undercharges by assuming fixed membership. In extensive-form games, determination depth measures the game’s strategic depth—the number of moves where a player must break a tie among equally-optimal options—which can be strictly less than the game-tree depth; one player can influence the other’s strategic depth, creating a new dimension of strategic interaction (Appendix D.2). In the distributed setting, the framework recovers the Halpern–Moses impossibility of asynchronous common knowledge [HM90], with the conservation law and exponential separation providing quantitative predictions beyond the binary threshold.
@@ 2 Framework
A commitment $\varphi$ is a constraint on outcomes. We write $H\cdot\varphi$ for the history obtained by adding a commitment event $\varphi$ after all maximal events of $H$ (i.e., $e\to\varphi$ for every $e$ with no successor in $H$), giving $H\sqsubseteq H\cdot\varphi$. Adding a commitment can only narrow the admissible set (shrinkage): $\mathsf{Spec}(H\cdot\varphi)\subseteq\mathsf{Spec}(H)$.
@@
Each commitment event $\varphi$ in a history carries its context: the prefix of the history up to $\varphi$ determines the admissible set at which $\varphi$ takes effect. Along any chain of extensions, environment events may enlarge or shrink the admissible set, but each commitment event can only shrink it (shrinkage condition). A relational specification cannot be resolved to a function by environment events alone; commitments are required, and the cost of those commitments is what determination depth measures (Section 2.1).
@@
The determination $D(H)$ of a history $H$ over a commitment basis $\Phi$ is the subsequence $\varphi_{1}\cdot\cdots\cdot\varphi_{m}$ of commitment events in $H$, listed in causal order. Its cost is $\mathsf{cost}(D(H))\triangleq m$.
The commitment events in any single-agent history are totally ordered (each follows all maximal events of its prefix), so the determination is a well-defined sequence. Writing $H_{j}$ for the prefix of $H$ through $\varphi_{j}$, shrinkage accumulates: $\mathsf{Spec}(H_{j})\subseteq\mathsf{Spec}(H_{j-1})$ for each $j$. A determination is valid if each $\varphi_{j}$ preserves nonemptiness: $\mathsf{Spec}(H^{\prime})\neq\emptyset$ for every $H^{\prime}\sqsupseteq H_{j}$. A specification is resolved if $|\mathsf{Spec}(H)|=1$ for all $H$; a history resolves $\mathsf{Spec}$ if its determination leaves $|\mathsf{Spec}(H)|=1$.
@@
Commitments $\varphi$ and $\psi$ commute with respect to $\mathsf{Spec}$ at $H$ if the order of extension does not matter: $\mathsf{Spec}(H\cdot\varphi\cdot\psi)=\mathsf{Spec}(H\cdot\psi\cdot\varphi)$. Commutation is relative to a specification and may change as the history grows: $\varphi$ and $\psi$ may commute at $H$ but not at an extension $H^{\prime}\sqsupseteq H$, or vice versa. This history-dependence does not arise in classical trace theory, where independence is static, and it is what makes non-commuting commitments impose irreducible sequencing. (A history records one linearization of the commitments; commutation ensures the choice of linearization is immaterial.) The depth of a determination is the total number of commuting layers across all runs of consecutive commitments, where commitments within each layer commute pairwise with respect to the specification at that point in the history. The full algebraic development appears in Section A.
@@
A determination strategy $\sigma$ resolves $\mathsf{Spec}$ if, for every history $H$ arising under $\sigma$, $|\mathsf{Spec}(H)|=1$ once all of $\sigma$’s commitments have been applied. Define $\mathsf{Cost}_{\Phi}(\mathsf{Spec})\triangleq\min_{\sigma}\;\max_{H}\;\mathsf{cost}(D(H))$ and $\mathsf{Depth}_{\Phi}(\mathsf{Spec})\triangleq\min_{\sigma}\;\max_{H}\;0pt(D(H))$, where $\sigma$ ranges over strategies that resolve $\mathsf{Spec}$ and $H$ ranges over histories arising under $\sigma$; or $\infty$ if no resolving strategy exists. In the offline setting (a single history with no environment extensions), the max is trivial and the definition reduces to a min over determinations.
$\mathsf{Cost}_{\Phi}(\mathsf{Spec})$ is the determination cost and $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ the determination depth. Both are properties of the specification, not of any particular strategy—the strategy is the witness, just as a circuit witnesses the depth of a function—and they are the semantic analogues of circuit size and circuit depth: cost counts commitments, depth counts sequential layers. In the online setting, the max over $H$ gives the environment adversarial control bounded by the history class $\mathcal{H}$; this is a worst-case measure, and the results in this paper focus on this setting.
@@ 2.1 Commitment bases
Atomic commitments are the finest-grained irrevocable commitments: each permanently excludes a single outcome at a single history. The permanence condition ($o$ never reappears at any extension) is what makes the exclusion irrevocable; it is satisfied by all specifications studied in this paper.
In online settings, depth under the atomic basis arises from forward validity constraints: two atomic commitments made at history $H$ cannot be applied in the same layer if their joint effect empties the admissible set at an extension of $H$, as in the following example:
@@
Let $O=\{a,b,c\}$ with $\mathsf{Spec}(H_{0})=\{a,b,c\}$, and suppose the environment may extend $H_{0}$ to a history $E\sqsupseteq H_{0}$ with $\mathsf{Spec}(E)=\{a,b\}$. The refined admissible sets $\mathsf{Spec}(H_{0}\cdot\varphi_{\neg a})$ and $\mathsf{Spec}(E\cdot\varphi_{\neg a})$ are both non-empty ($\{b,c\}$ and $\{b\}$ respectively), so $\varphi_{\neg a}$ is valid at $H_{0}$. But the joint refinement $\mathsf{Spec}(E\cdot\varphi_{\neg a}\cdot\varphi_{\neg b})$ is empty: $\{a,b\}\setminus\{a,b\}=\emptyset$. Symmetrically, if the environment may also extend $H_{0}$ to $E^{\prime}$ with $\mathsf{Spec}(E^{\prime})=\{b,c\}$ and to $E^{\prime\prime}$ with $\mathsf{Spec}(E^{\prime\prime})=\{a,c\}$, then no pair of atomic commitments at $H_{0}$ is jointly valid, so resolution requires two layers: commit, observe the environment, then commit again.
@@
Intrinsic depth measures the irreducible layering forced by forward validity constraints alone—the “bare” sequential cost before any computational or architectural assumptions are introduced. In the offline setting, forward validity constraints vanish and intrinsic depth collapses; the conservation law (Theorem 3) shows that the sequential cost reappears as computation within layers.
@@
A commitment $\varphi$ is pointwise if it acts as an outcome-by-outcome filter: whether $o\in\mathsf{Spec}(H\cdot\varphi)$ depends only on $o$ and $H$, not on the rest of $\mathsf{Spec}(H)$. A pointwise basis consists entirely of pointwise commitments. A commitment whose effect on one outcome depends on the presence of another is non-pointwise.
@@
The atomic basis is both commutative and constant-depth, making it the canonical minimal basis for the online setting where commitments are made before all information has arrived. Under a constant-depth basis, every unit of determination depth corresponds to one unit of sequential time; under a richer basis, some of that time is absorbed into the computation needed to evaluate each commitment. The pointwise, commutative, and constant-depth properties each appear in multiple results: pointwise bases arise in the exponential separation (Section 3), the PH characterization (Section 5), and the NP-hardness of computing depth; commutative bases in the oracle characterization (Section 4.2) and the distributed extension (Appendix B); and constant-depth bases in the conservation law (Section 4.1).
@@ 3 Exponential Depth–Width Separation
We exhibit a family of constrained generation tasks—modeling autoregressive text generation—in which every commitment is a constant-time table lookup, yet any strategy using $d^{\prime}<k$ layers requires width at least $(m/s)^{k-d^{\prime}}$.
@@ 3.1 The constrained generation task
Fix positive integers $k$ (number of positions), $m$ (domain size), and $s$ (constraint sparsity), with $1\leq s\leq m$. The task has k1k-1 links $1\to 2\to\cdots\to k$, each constraining a position’s value given its predecessor’s.
@@
The specification is relational ($s^{k}$ admissible tuples when $s\geq 2$) and models autoregressive generation: each position is a token, $[m]$ is the vocabulary, and $P_{\ell}$ encodes local coherence constraints.
@@
Committing to one position per layer in order $v_{1},v_{2},\ldots,v_{k}$ always produces a valid tuple (each $v_{\ell}\in P_{\ell}(v_{\ell-1},\cdot)$ exists by sparsity), using $k$ layers and width $1$.
@@ 3.2 Strategy Model and Lower Bound
Definition 14 ($d^{\prime}$-layer strategy with width $w$).
A $d^{\prime}$-layer strategy with width $w$ partitions the $k$ positions into $d^{\prime}$ groups $S_{1},\ldots,S_{d^{\prime}}$ and produces a single history carrying $w$ parallel candidate tuples, each contributing $k$ commitments (one per position) distributed across the $d^{\prime}$ layers. In layer $r$, for each candidate, the strategy commits to values at all positions in $S_{r}$; these values may depend on the full input and on earlier-layer commitments, but not on same-layer or later-layer commitments. The strategy selects one of the $w$ candidates as its output; it resolves the specification if the output lies in $\mathsf{Spec}_{\mathsf{cg}}$.
The model is strictly more permissive than autoregressive generation (which fixes left-to-right order and width $1$): it allows arbitrary layer assignments, arbitrary within-layer parallelism, and $w$ independent candidates. The only constraint is that commitments within a layer cannot depend on same-layer outcomes—the defining property of parallel execution—so any lower bound here applies to every parallel strategy that respects this constraint, including beam search, speculative decoding, and diffusion-style refinement.
Each group $S_{r}$ is a commuting layer in which the $w$ candidates each commit to $|S_{r}|$ positions in parallel. When $d^{\prime}<k$, some adjacent positions share a layer and cannot observe each other’s values; the lower bound counts these blind spots (uninformed links) and shows that $w$ grows exponentially in their number.
@@
I.e., membership of any single element in a row has probability $\leq\gamma$, conditioned on the rest of the chain.
@@
Fix a single candidate tuple $(v_{1},\ldots,v_{k})$ and process the uninformed links in order. At link $\ell_{j}$, condition on all constraint data except the row $P_{\ell_{j}}(v_{\ell_{j}-1}^{*},\cdot)$ and on the outcomes at prior links. Under this conditioning, both $v_{\ell_{j}-1}^{*}$ (from an earlier layer or the conditioning) and $v_{\ell_{j}}$ (from the same or a later layer) are fixed, while the row $P_{\ell_{j}}(v_{\ell_{j}-1}^{*},\cdot)$ remains random by the $\gamma$-spread condition. Hence $\Pr[(v_{\ell_{j}-1}^{*},v_{\ell_{j}})\in P_{\ell_{j}}\mid\text{conditioning}]\leq\gamma$. Multiplying across all $t\geq k-d^{\prime}$ links gives success probability at most $\gamma^{k-d^{\prime}}$ per candidate; a union bound over $w$ candidates gives the result. ∎
@@
Theorem 1(b) bounds the expected resolution probability of any deterministic strategy over the random $(k,m,s)$-distribution by $w\cdot(s/m)^{k-d^{\prime}}$, so for each deterministic strategy some chain achieves this bound. Yao’s minimax principle [YAO77] gives the same guarantee for randomized strategies.
Orthogonality.
Determination depth and computational depth are independent axes: the constrained generation task has determination depth $k$ and computational depth $O(1)$ per step, while any computationally hard function has determination depth 0. Within a single problem, however, the two axes interact: Section 4.1 shows that enriching the basis trades determination layers for computational depth within layers, but the total sequential cost is bounded below (Theorem 3).
@@
When a strategy tolerates constraint violations, the expected violation count is at least $(k-d^{\prime})(1-\gamma)$ by Theorem 1; for the random $(k,m,s)$-distribution this is tight. The parameter $\gamma$ thus controls the exchange rate between distributional knowledge and sequential depth. For a distribution that is not $\gamma$-spread—one with exploitable correlations—a strategy can achieve per-link success probability $\gamma^{\prime}>\gamma$. At $\gamma^{\prime}=1$ (perfect prediction), no layers are needed; at $\gamma^{\prime}=\gamma$ (no exploitable structure), the full $k$ layers are required.

### 01 Apr 2026 v2

@@ Abstract
Classical complexity theory measures the cost of computing a function, but many computational tasks require committing to one valid output among several. We introduce determination depth—the minimum number of sequential layers of irrevocable commitments needed to select a single valid output—and show that no amount of computation can eliminate this cost. We exhibit relational tasks whose commitments are constant-time table lookups yet require exponential parallel width to compensate for any reduction in depth. A conservation law shows that enriching commitments merely relabels determination layers as circuit depth, preserving the total sequential cost. For circuit-encoded specifications, the resulting depth hierarchy captures the polynomial hierarchy ($\Sigma_{2k}^{P}$-complete for each fixed $k$, PSPACE-complete for unbounded $k$). In the online setting, determination depth is fully irreducible: unlimited computation between commitment layers cannot reduce their number.
@@ 1 Introduction
Classical complexity theory measures the cost of computing a function—producing the unique correct output. But many tasks require resolving a relation: choosing one admissible outcome from a set, as when a text generator selects one coherent continuation among many, or a distributed system must agree on one valid output among several. We show that resolving a relation carries an intrinsic sequential cost that no amount of computation can eliminate. Classical hard functions are computationally expensive but have no cost on this axis. However, there exist relational tasks in which every step is a constant-time table lookup, yet any strategy using fewer than $k$ sequential layers requires exponential width, even with unbounded parallelism within each layer. The bottleneck is not computation, nor communication, but irrevocable commitment. This cost is invisible to classical complexity measures—circuit depth, communication complexity, adaptive queries—which track the cost of computing or distributing an answer, not the cost of committing to one.
@@
We establish three main results. (1) Exponential separation (Section 3): the depth–width tradeoff is exponential, generalizing pointer chasing [NW93, MYZ25] from functions to relations. (2) Computation cannot help (Section 4): the sequential cost of determination is governed by the dependency structure among commitments. Enriching the computational power of commitments merely relabels determination layers as circuit depth within layers—the total sequential cost is conserved (conservation law); the bound is tight. Enriching the computational power of the strategy does not help either: under a commutative basis in the online setting, an oracle characterization shows that determination depth equals the minimum number of observations of the growing history, regardless of computation between observations. (3) Polynomial-hierarchy characterization (Section 5): for circuit-encoded specifications, “is depth $\leq k$?” is $\Sigma_{2k}^{P}$-complete for each fixed $k$ and PSPACE-complete for unbounded $k$, certifying determination depth as a semantic measure well-calibrated against the classical hierarchy.
@@
The appendices exhibit both phenomena. BSP round complexity overcharges by counting communication rounds that involve no commitment, and undercharges by assuming fixed membership. In extensive-form games, determination depth bounds the game’s strategic depth—the number of moves where a player must break a tie among equally-optimal options—which can be strictly less than the game-tree depth; one player can influence the other’s strategic depth, creating a new dimension of strategic interaction (Appendix D.2). In the distributed setting, the framework recovers the Halpern–Moses impossibility of asynchronous common knowledge [HM90], with the conservation law and exponential separation providing quantitative bounds beyond the binary threshold.
@@ 2 Framework
A commitment $\varphi$ is a constraint on outcomes. We write $H\cdot\varphi$ for the history obtained by adding a commitment event $\varphi$ after all maximal events of $H$ (i.e., $e\to\varphi$ for every $e$ with no successor in $H$), giving $H\sqsubseteq H\cdot\varphi$. Adding a commitment permanently narrows the admissible set (shrinkage): $\mathsf{Spec}(H^{\prime})\subseteq\mathsf{Spec}(H)$ for every $H^{\prime}\sqsupseteq H\cdot\varphi$.
@@
Each commitment event $\varphi$ in a history carries its context: the prefix of the history up to $\varphi$ determines the admissible set at which $\varphi$ takes effect. Environment events may change the admissible set in either direction, but commitments can only shrink it, permanently (shrinkage condition). Not every relational specification requires commitments; quantifying the requirement is the theme of this work.
@@
The determination $D(H)$ of a history $H$ over a commitment basis $\Phi$ is the sequence $\varphi_{1}\cdot\cdots\cdot\varphi_{m}$ ($\varphi_{i}\in\Phi$) of all commitment events in $H$, listed in causal order. Its cost is $\mathsf{cost}(D(H))\triangleq m$.
The commitment events in any single-agent history are totally ordered (each follows all maximal events of its prefix), so the determination is a well-defined sequence. Writing $H_{j}$ for the prefix of $H$ through $\varphi_{j}$, shrinkage accumulates: $\mathsf{Spec}(H_{j})\subseteq\mathsf{Spec}(H_{j-1})$ for each $j$. A determination is valid if each $\varphi_{j}$ preserves nonemptiness: $\mathsf{Spec}(H^{\prime})\neq\emptyset$ for every $H^{\prime}\sqsupseteq H_{j}$. The goal is to narrow the admissible set to a single outcome ($|\mathsf{Spec}(H)|=1$); the formal notion of a resolving strategy appears in Definition 6.
@@
Example 1 (Consensus server).
Consensus—agreeing on a single value—is a canonical coordination problem in distributed systems. A single sequential server that dictates a choice for the system is a simplified implementation, but sufficient to illustrate the three forces. The consensus specification maps a history to a set of candidate values. Proposals arrive over time (environment events, extending history $H\sqsubseteq H^{\prime}$), each adding a value to the admissible set at the extended history. The strategy’s determination consists of two commitments. First, $\varphi_{c}$ closes the proposal phase at some history $H_{c}$: after $H_{c}\cdot\varphi_{c}$, the specification ignores subsequent proposals, so no environment extension of $H_{c}\cdot\varphi_{c}$ can enlarge the admissible set. Then $\varphi_{w}$ draws a winner, resolving the specification to one value from the closed proposal set. The draw depends on the closure—the strategy cannot choose a winner without knowing the candidates—so the two commitments require two sequential layers.
@@
Commitments $\varphi$ and $\psi$ commute with respect to $\mathsf{Spec}$ at $H$ if the order of extension does not matter: $\mathsf{Spec}(H\cdot\varphi\cdot\psi)=\mathsf{Spec}(H\cdot\psi\cdot\varphi)$. Commutation is relative to a specification and may change as the history grows: $\varphi$ and $\psi$ may commute at $H$ but not at an extension $H^{\prime}\sqsupseteq H$, or vice versa. This history-dependence does not arise in classical trace theory [MAZ77], where independence is static; in our setting, non-commuting commitments impose irreducible sequencing precisely because their commutation status can shift as the history evolves. (A history records one linearization of the commitments; commutation ensures the choice of linearization is immaterial.) The depth of a determination is the minimum number of commuting layers needed to organize its commitments: within each layer, all commitments commute pairwise; across layers, they need not. In the online setting, environment events may separate consecutive commitments into distinct runs, and the depth is the total number of layers across all runs (formalized in Section A).
@@
A determination strategy $\sigma$ resolves $\mathsf{Spec}$ if, for every history $H$ arising under $\sigma$, $|\mathsf{Spec}(H)|=1$ once all of $\sigma$’s commitments have been applied. Define $\mathsf{Cost}_{\Phi}(\mathsf{Spec})\triangleq\min_{\sigma}\;\max_{H}\;\mathsf{cost}(D(H))$ and $\mathsf{Depth}_{\Phi}(\mathsf{Spec})\triangleq\min_{\sigma}\;\max_{H}\;0pt(D(H))$, where $\sigma$ ranges over strategies that resolve $\mathsf{Spec}$ and $H$ ranges over histories arising under $\sigma$; or $\infty$ if no resolving strategy exists. In the offline setting (a single complete history), a strategy is simply a determination of that history, and the definition reduces to a min over determinations.
$\mathsf{Cost}_{\Phi}(\mathsf{Spec})$ is the determination cost and $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ the determination depth. Both are properties of the specification, not of any particular strategy—the strategy is the witness, just as a circuit witnesses the depth of a function. Cost counts commitments; depth counts sequential layers. In the online setting, the max over $H$ gives the environment adversarial control bounded by the history class $\mathcal{H}$; this is a worst-case measure, and the results in this paper focus on this setting. In the consensus server example (Example 1), the determination cost is $2$ (close proposals, then draw) and the determination depth is also $2$: the draw depends on the closure, so the two commitments cannot share a layer.
@@ 2.1 Commitment bases
Atomic commitments are the finest-grained irrevocable commitments: each permanently excludes a single outcome at a single history.
In online settings, depth under the atomic basis arises from forward validity constraints: two atomic commitments made at history $H$ cannot be applied in the same layer if their joint effect empties the admissible set at an extension of $H$, as in the following consensus example. In this variant, environment events may also shrink the admissible set—e.g., by retracting proposals:
@@
Consider a consensus instance (cf. Example 1) with three candidate values $O=\{a,b,c\}$ and $\mathsf{Spec}(H_{0})=\{a,b,c\}$. Suppose the environment may extend $H_{0}$ to a history $E\sqsupseteq H_{0}$ with $\mathsf{Spec}(E)=\{a,b\}$. The refined admissible sets $\mathsf{Spec}(H_{0}\cdot\varphi_{\neg a})$ and $\mathsf{Spec}(E\cdot\varphi_{\neg a})$ are both non-empty ($\{b,c\}$ and $\{b\}$ respectively), so $\varphi_{\neg a}$ is valid at $H_{0}$. But the joint refinement $\mathsf{Spec}(E\cdot\varphi_{\neg a}\cdot\varphi_{\neg b})$ is empty: $\{a,b\}\setminus\{a,b\}=\emptyset$. Symmetrically, if the environment may also extend $H_{0}$ to $E^{\prime}$ with $\mathsf{Spec}(E^{\prime})=\{b,c\}$ and to $E^{\prime\prime}$ with $\mathsf{Spec}(E^{\prime\prime})=\{a,c\}$, then no pair of atomic commitments at $H_{0}$ is jointly valid, so resolution requires two layers: commit, observe the environment, then commit again.
@@
Intrinsic depth measures the irreducible sequential cost of resolving a specification under the finest-grained commitments, before any computational or architectural assumptions are introduced.
@@
A commitment $\varphi$ is pointwise if it acts as an independent filter on each outcome: whether $o\in\mathsf{Spec}(H\cdot\varphi)$ depends only on $o$ and $H$, not on which other outcomes belong to $\mathsf{Spec}(H)$. A pointwise basis consists entirely of pointwise commitments.
@@
The atomic basis is both commutative and constant-depth, making it the canonical minimal basis for the online setting where commitments are made before all information has arrived. The pointwise, commutative, and constant-depth properties each appear in the results that follow.
@@ 3 Exponential Depth–Width Separation
This section proves that determination depth is a meaningful barrier: reducing it requires an exponential blowup in parallel resources, even when each individual commitment is trivial to compute. The consensus server (Example 1) required two sequential layers; we now exhibit tasks requiring $k$ layers for any $k$. We construct a family of constrained generation tasks—modeling autoregressive text generation—in which every commitment is a constant-time table lookup, yet any strategy using $d^{\prime}<k$ layers requires width at least $(m/s)^{k-d^{\prime}}$.
@@ 3.1 The constrained generation task
Fix positive integers $k$ (number of positions), $m$ (domain size), and $s$ (constraint sparsity), with $1\leq s\leq m$.
@@
The specification is relational ($s^{k}$ admissible tuples when $s\geq 2$) and models autoregressive generation: each position is a token, $[m]$ is the vocabulary, and the k1k-1 links $1\to 2\to\cdots\to k$ encode local coherence constraints.
@@
Committing to one position per layer in order $v_{1},v_{2},\ldots,v_{k}$ always produces a valid tuple—each $v_{\ell}\in P_{\ell}(v_{\ell-1},\cdot)$ exists by sparsity—using $k$ layers and width $1$.
@@ 3.2 Strategy Model and Lower Bound
Definition 15 ($d^{\prime}$-layer strategy with width $w$).
A $d^{\prime}$-layer strategy with width $w$ partitions the $k$ positions into $d^{\prime}$ groups $S_{1},\ldots,S_{d^{\prime}}$ and builds $w$ candidate tuples in parallel. In layer $r$, the strategy assigns values to the positions in group $S_{r}$ for each candidate, depending only on the full input and earlier-layer assignments. The strategy resolves the specification if any candidate lies in $\mathsf{Spec}_{\mathsf{cg}}$; the final selection among valid candidates adds at most one layer, which does not affect the asymptotic lower bound.
At each uninformed link, the strategy must guess a value without knowing its predecessor; the theorem below shows that each guess fails with probability controlled by a distributional parameter $\gamma$.
@@
In short, membership of any one element in a row has probability $\leq\gamma$, conditioned on the rest of the chain.
@@
The model is strictly more permissive than autoregressive generation (which fixes left-to-right order and width $1$): it allows arbitrary layer assignments, arbitrary within-layer parallelism, and $w$ independent candidates. The only constraint is that commitments within a layer cannot depend on same-layer outcomes—the defining property of parallel execution—so the lower bound applies to every parallel strategy that respects this constraint, including beam search, speculative decoding, and diffusion-style refinement.
@@
Fix a single candidate tuple $(v_{1},\ldots,v_{k})$ and consider an uninformed link $\ell_{j}$. The strategy picks $v_{\ell_{j}}$ without seeing $P_{\ell_{j}}(v_{\ell_{j}-1}^{*},\cdot)$, which is still $\gamma$-spread by assumption, so $\Pr[(v_{\ell_{j}-1}^{*},v_{\ell_{j}})\in P_{\ell_{j}}]\leq\gamma$. Multiplying across all $t\geq k-d^{\prime}$ links gives success probability at most $\gamma^{k-d^{\prime}}$ per candidate; a union bound over $w$ candidates gives the result. ∎
@@
Theorem 1(b) bounds the expected resolution probability of any deterministic strategy over the random $(k,m,s)$-distribution by $w\cdot(s/m)^{k-d^{\prime}}$, so for each deterministic strategy some chain exhibits this bound. Yao’s minimax principle [YAO77] gives the same guarantee for randomized strategies.
@@
When a strategy tolerates constraint violations, the expected violation count is at least $(k-d^{\prime})(1-\gamma)$ by Theorem 1; for the random $(k,m,s)$-distribution this is tight. The spread parameter $\gamma$ controls the exchange rate between distributional structure and sequential depth: a distribution with more exploitable structure has a smaller effective $\gamma$, reducing the layers needed. At one extreme ($\gamma\to 0$, perfect prediction), no layers are needed; at the other ($\gamma=s/m$, no exploitable structure), the full $k$ layers are required.